# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v3

@@ 5.8 Cauchy–Pompeiu Representation Theorem
Chapter 6 Covariant Holomorphicity and Gauge Structure
Covariant Holomorphicity and Gauge Structure in the Rigid Regime
@@ 8.1 The Rigid Variable Laplace Operator
In the present framework, once rigidity is imposed Chapter  2, $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (5) becomes available. It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
@@ 9.1.1 Elliptic system and induced structure
and $\alpha,\beta$ determine a variable elliptic algebra
@@ 9.1.4 Elliptic conjugation and Vekua form
Using
$U=\tfrac{1}{2}(W+\widehat{W}),\qquad V=\frac{W-\widehat{W}}{2i+\beta},$
the right–hand side of (9.5) can be written uniquely as
$F-AW-B\widehat{W}$,
with
$F:=\tfrac{1}{2}(f+ig),$
and coefficients
$A$ $=A_{0}-\frac{1}{2}\,\frac{G}{2i+\beta},$
$B$ $=B_{0}+\frac{1}{2}\,\frac{G}{2i+\beta},$
@@
$A_{0}$ $=\frac{1}{4}\Big(a-\frac{\beta}{\alpha}b+d+i\,c-i\,\frac{1}{\alpha}b\Big),$
$B_{0}$ $=\frac{1}{4}\Big(a+\frac{\beta}{\alpha}b-d+i\,c+i\,\frac{1}{\alpha}b\Big).$
Dividing (9.5) by $2$, we arrive at the canonical variable–structure Vekua equation
@@ 9.1.5 Remarks
  • When the structure is rigid ($G\equiv 0$), the coefficients reduce to their classical forms.
  • The obstruction $G$ measures precisely the deviation from rigid Cauchy–Riemann calculus and is absorbed into the lower–order terms.
@@ 9.2 Rigidization
$\alpha,\beta\in C^{1}(\Omega),\qquad\Delta=4\alpha-\beta^{2}>0.$
@@ 9.2.2 Transport form of the obstruction
$G\equiv 0\quad\Longleftrightarrow\quad\lambda_{x}+\lambda\,\lambda_{y}=0.$
@@ 9.2.3 Definition of rigidization
Rigidization is an Eulerian construction: the diffeomorphism is generated by an ODE driven by the transport field of the structure.
9.2.4 Rigidization ODE
Let $\lambda(x,y)$ be the spectral slope of the original structure. Define the vector field
$V:=\partial_{x}+\lambda\,\partial_{y}.$
Proposition 9.4 (Rigidizing flow).
Let $(X(x,y),Y(x,y))$ solve the first–order system
$\begin{cases}V(X)=1,\\ V(Y)=\lambda.\end{cases}$ (9.8)
Then $(X,Y)$ defines a local diffeomorphism, and in the (X,Y)(X,Y)–coordinates the transport equation becomes
$\lambda_{X}+\lambda\,\lambda_{Y}=0.$
@@ 9.2.4 Rigidization ODE
Along the flow of $V$, equation (9.8) ensures that $X$ is the flow parameter and $Y$ is transported consistently with $\lambda$. By construction, derivatives along $V$ become $\partial_{X}$, eliminating the source term in (9.7). ∎
9.2.5 Algorithmic description
Rigidization can be implemented as the following local algorithm.
@@ 9.2.5 Algorithmic description
    Compute $\alpha,\beta$ from the principal part of the system.
@@
    Compute $\Delta=4\alpha-\beta^{2}$ and the spectral slope
    $\lambda=\frac{-\beta+i\sqrt{\Delta}}{2}.$
@@
    Form the transport vector field
    $V=\partial_{x}+\lambda\,\partial_{y}.$
@@
    Solve the first–order ODE system
    $V(X)=1,\qquad V(Y)=\lambda$
    with prescribed initial conditions.
@@
    Express the system in (X,Y)(X,Y)–coordinates and recompute the induced structure polynomial.
The resulting structure is rigid by construction.
@@ 9.2.7 Remarks
  • Rigidization requires solving only first–order ODEs; no Beltrami equation or singular integral is involved.
@@
  • Rigidization separates geometric transport from analytic solving: once rigid, the system admits the full rigid Cauchy–Vekua calculus.

### v4

@@ 5.8 Cauchy–Pompeiu Representation Theorem
Chapter 6 Covariant Holomorphicity and Gauge Structure in the Rigid Regime
@@ 8.1 The Rigid Variable Laplace Operator
In the present framework, once rigidity is imposed (Chapter 2), $\partial_{\bar{z}}$ becomes a derivation and the Cauchy–Pompeiu formula (Chapter 5) becomes available. It is then natural—and unavoidable—to study the second–order operator obtained by composing $\partial_{\bar{z}}$ with its conjugate partner $\partial_{z}$.
@@ 9.1.1 Elliptic system and induced structure
and $\alpha,\beta$ determine a variable elliptic algebra bundle
@@
For an algebra–valued field $W=U+iV$ with $U,V$ real, we write
$\widehat{W}:=U+\widehat{i}\,V=U-(\beta+i)V.$
@@ 9.1.4 Elliptic conjugation and Vekua form
The conjugate field $\widehat{W}$ satisfies the linear identities
$W-\widehat{W}=(2i+\beta)V,\qquad W+\widehat{W}=2U-\beta V.$ (9.6)
Since $\Delta>0$, the element $2i+\beta$ is invertible in $A_{z}$ and in fact
$(2i+\beta)^{-1}=\frac{\beta-2i}{\beta^{2}+4\alpha}.$
Therefore we may solve (9.6) for $U$ and $V$ in terms of (W,W^)($\widehat{W}$):
$V=\frac{W-\widehat{W}}{2i+\beta},\qquad U=\frac{1}{2}\left(W+\widehat{W}+\beta\,\frac{W-\widehat{W}}{2i+\beta}\right).$ (9.7)
Substituting (9.7) into the right–hand side of (9.5) expresses it as an $A_{z}$–linear combination of $W$ and $\widehat{W}$. Concretely, there exist unique algebra–valued coefficients $A_{0},B_{0}$ and a unique datum $F_{0}$ such that
$(f+ig)-(a+ic)U-(b+id)V\;=\;2F_{0}-2A_{0}\,W-2B_{0}\,\widehat{W}$. (9.8)
(Existence and uniqueness follow because (9.7) is an invertible linear change of variables between $(U,V)$ and (W,W^)($\widehat{W}$).)
Define
$F:=F_{0}=\tfrac{1}{2}(f+ig),$
and absorb the obstruction term $VG$ from (9.5) using $V=(W-\widehat{W})/(2i+\beta)$:
$VG=\frac{W-\widehat{W}}{2i+\beta}\,G.$
Thus (9.5) may be rewritten uniquely in the form
$2\bar{\partial}W=2F-2A\,W-2B\,\widehat{W}$, (9.9)
@@
$A\;:=\;A_{0}-\frac{1}{2}\,\frac{G}{2i+\beta},\qquad B\;:=\;B_{0}+\frac{1}{2}\,\frac{G}{2i+\beta}.$ (9.10)
Dividing by $2$ yields the canonical variable–structure Vekua equation
@@ 9.1.5 Remarks
  • When the structure is rigid ($G\equiv 0$), the obstruction contribution in (9.10) vanishes and the lower–order coefficients reduce to the rigid forms determined by (9.8).
  • The obstruction $G$ measures precisely the deviation from rigid Cauchy–Riemann calculus and is absorbed into the lower–order terms via the explicit splitting in (9.10).
@@ 9.2 Rigidization
$\alpha,\beta\in C^{2}(\Omega),\qquad\Delta=4\alpha-\beta^{2}>0.$
@@ 9.2.2 Transport form of the obstruction
$G\equiv 0\;\Longleftrightarrow\;\lambda_{x}+\lambda\,\lambda_{y}=0.$
@@ 9.2.3 Definition of rigidization
9.2.4 The characteristic operator and its square
The key analytic object is the characteristic operator
$V:=\partial_{x}+\lambda\,\partial_{y},$
where $\lambda=a+ib$ with $a=\operatorname{Re}\lambda$ and $b=\operatorname{Im}\lambda>0$. Under a real change of coordinates $(x,y)\mapsto(X,Y)$, the operator $V$ acts on the new coordinate functions by
$V(X)=X_{x}+\lambda\,X_{y},\qquad V(Y)=Y_{x}+\lambda\,Y_{y}.$
The transformed spectral slope is
$\widetilde{\lambda}\;=\;\frac{V(Y)}{V(X)},$ (9.13)
and the rigidization condition $\widetilde{\lambda}_{X}+\widetilde{\lambda}\,\widetilde{\lambda}_{Y}=0$ is equivalent to the ratio condition
$V\!\bigl(\widetilde{\lambda}\bigr)=0\qquad\Longleftrightarrow\qquad V^{2}(Y)\cdot V(X)\;=\;V(Y)\cdot V^{2}(X).$ (9.14)
The iterated operator $V^{2}=(\partial_{x}+\lambda\,\partial_{y})^{2}$ admits a canonical decomposition when applied to a real–valued function $\varphi\in C^{2}(\Omega,\mathbb{R})$. A direct computation using $\lambda=a+ib$ yields
$V^{2}(\varphi)\;=\;W(\varphi)\;+\;i\,T(\varphi),$ (9.15)
where $T$ and $W$ are the real operators defined below.
Proposition 9.5 ($V^{2}$ decomposition).
Let $\varphi\in C^{2}(\Omega,\mathbb{R})$ and write $H=H_{R}+iH_{I}$ for the forcing (9.12). Then
$T(\varphi)$ $=2b\Bigl[(\varphi_{y})_{x}+a\,(\varphi_{y})_{y}+\frac{H_{I}}{2b}\,\varphi_{y}\Bigr],$ (9.16)
$W(\varphi)$ $=\varphi_{xx}-|\lambda|^{2}\,\varphi_{yy}+\Bigl(H_{R}-\frac{a\,H_{I}}{b}\Bigr)\varphi_{y}.$ (9.17)
@@ 9.2.4 The characteristic operator and its square
A direct expansion gives
$V^{2}(\varphi)=\varphi_{xx}+2\lambda\,\varphi_{xy}+\lambda^{2}\,\varphi_{yy}+(\lambda_{x}+\lambda\,\lambda_{y})\,\varphi_{y}.$
The imaginary part of this expression, using $\lambda^{2}=(a^{2}-b^{2})+2abi$ and $\lambda_{x}+\lambda\,\lambda_{y}=H_{R}+iH_{I}$, is
$\operatorname{Im}V^{2}(\varphi)=2b\,\varphi_{xy}+2ab\,\varphi_{yy}+H_{I}\,\varphi_{y},$
which coincides with (9.16). Setting this equal to zero and eliminating $\varphi_{xy}$ from the real part yields (9.17). ∎
9.2.5 Transport and wave structure
The two components in Proposition 9.5 have distinct and complementary PDE character.
Transport component. Setting $p:=\varphi_{y}$, equation $T(\varphi)=0$ becomes the first–order linear transport equation
$p_{x}+a\,p_{y}=-\frac{H_{I}}{2b}\,p,$ (9.18)
whose transport direction is the real vector field $(1,\,\operatorname{Re}\lambda)$. This equation is solvable by the method of real characteristics: along the curves $\tfrac{dy}{dx}=a(x,y)$, the function $p$ satisfies a linear ODE.
Wave component. The equation $W(\varphi)=0$ is the real second–order equation
$\varphi_{xx}-|\lambda|^{2}\,\varphi_{yy}+c\,\varphi_{y}=0,\qquad c:=H_{R}-\frac{a\,H_{I}}{b},$ (9.19)
whose principal symbol $\xi^{2}-|\lambda|^{2}\eta^{2}=({\xi-|\lambda|\eta})({\xi+|\lambda|\eta})$ is hyperbolic, with real characteristic speeds $\pm|\lambda|^{-1}$.
Neither component alone is elliptic. The ellipticity of the full rigidization problem arises from their coupling, in exact analogy with the Cauchy–Riemann equations $(u_{x}=v_{y},\;u_{y}=-v_{x})$, which are elliptic despite each equation individually describing simple transport.
9.2.6 PDE type of the rigidization problem
Theorem 9.6 (PDE type of rigidization).
Let $\lambda=a+ib$ with $b>0$ satisfy $\lambda_{x}+\lambda\,\lambda_{y}=H\not\equiv 0$. The ratio condition (9.14), viewed as a second–order system for two real unknown functions $(X,Y)$, is elliptic.
Proof.
The principal symbol of the ratio condition inherits its type from $V^{2}$. The characteristic equation $(\xi+\lambda\eta)^{2}=0$ has solutions $\xi/\eta=-\lambda=-a-ib$, which are complex since $b>0$. Hence no real characteristic directions exist, and the system is elliptic. ∎
Proposition 9.7 (Transport–transport incompatibility).
If $T(X)=0$ and $T(Y)=0$ simultaneously, then the ratio condition (9.14) forces the Jacobian $J=X_{x}Y_{y}-X_{y}Y_{x}$ to vanish.
Proof.
Under $T(X)=T(Y)=0$, the values $V^{2}(X)=W(X)$ and $V^{2}(Y)=W(Y)$ are real. Since $V(X)$ and $V(Y)$ are complex, the imaginary part of (9.14) reads
$W(Y)\cdot b\,X_{y}=W(X)\cdot b\,Y_{y}.$
The real part similarly gives $W(Y)(X_{x}+aX_{y})=W(X)(Y_{x}+aY_{y})$. Cross–multiplying the two proportionalities yields $X_{y}Y_{x}=Y_{y}X_{x}$, hence $J=0$. ∎
Remark 9.8.
Proposition 9.7 shows that the transport and wave components cannot be decoupled for both coordinate functions simultaneously while preserving non–degeneracy. The elliptic coupling between $T$ and $W$ is load–bearing.
9.2.7 Splitting method for rigidization
Although the rigidization problem is elliptic, the decomposition $V^{2}=W+iT$ of Proposition 9.5 enables an alternating–direction iterative method in which each individual step uses only real characteristics.
Algorithm (Transport–wave splitting).
  1. 0.
    Initialize $X^{0}=x$, $Y^{0}=y$ (identity map).
@@ 9.2.7 Splitting method for rigidization
    Compute the forcing residual $H^{n}:=\widetilde{\lambda}^{\,n}_{X}+\widetilde{\lambda}^{\,n}\,\widetilde{\lambda}^{\,n}_{Y}$ in the current coordinates.
@@
    Transport half–step. Correct $Y^{n}$ by solving the first–order transport equation (9.18) for $\delta Y_{y}$ along the real characteristics of $(1,\operatorname{Re}\lambda)$, targeting $\operatorname{Im}H^{n}\to 0$.
@@
    Wave half–step. Correct $X^{n}$ by solving the hyperbolic wave equation (9.19) for $\delta X$ along the wave characteristics $\pm|\lambda|^{-1}$, targeting $\operatorname{Re}H^{n}\to 0$.
@@
    Update: $X^{n+1}=X^{n}+\delta X$, $Y^{n+1}=Y^{n}+\delta Y$.
@@
    If $|H^{n+1}|$ is below the desired tolerance, stop; otherwise return to step 1.
This procedure treats the elliptic coupling iteratively while solving only hyperbolic subproblems at each step. It is directly analogous to alternating–direction implicit (ADI) methods for elliptic equations and to Uzawa iteration for saddle–point systems.
9.2.8 Relationship to Beltrami uniformization
It is instructive to compare rigidization with full Beltrami uniformization. The latter seeks a diffeomorphism $\Phi$ such that $\Phi^{*}\lambda$ is constant, requiring the solution of the Beltrami equation $w_{\bar{z}}=\mu\,w_{z}$, an elliptic PDE whose general theory relies on singular integral operators (the Beurling–Ahlfors transform).
Rigidization is strictly weaker: it asks only that $H\to 0$, while $\lambda$ is permitted to remain nonconstant. Both problems are elliptic, but the rigidization problem inherits additional structure from the Burgers self–transport (the characteristic speed equals the transported quantity), which constrains the effective Beltrami coefficient and enables the splitting method described above.
@@ 9.2.10 Remarks
  • Rigidization does not reduce to solving first–order ODEs along characteristics. The ratio condition (9.14) is a genuinely elliptic problem for the coordinate functions $(X,Y)$. However, the transport–wave decomposition of $V^{2}$ provides a natural splitting in which each iterative half–step uses only real characteristics.
@@
  • Rigidization separates geometric transport from analytic solving: once rigid, the system admits the full rigid Cauchy–Vekua calculus (Chapters 5–8).
9.2.11 Geometric interpretation
The decomposition $V^{2}=W+iT$ expresses, at the PDE level, the central theme of this monograph. The transport component $T$ acts on $\varphi_{y}$—the transverse gradient—and propagates along the real projection $(1,\operatorname{Re}\lambda)$ of the complex characteristic direction. It belongs to the structural layer of Chapter 2. The wave component $W$ acts on $\varphi$ itself and propagates along the characteristic speeds $\pm|\lambda|^{-1}$ of the rigid elliptic operator $L_{\alpha,\beta}$ of Chapter 8. It belongs to the analytic layer.
Transport and integrability are therefore the two halves of the rigidization problem. Individually, each is hyperbolic; jointly, they are elliptic. Rigidization succeeds precisely when these two components can be made simultaneously compatible—the analytic expression of the passage from “transport first” to “integrability second” that governs the entire theory.